How to write an algebraic model and solve the problem.
You and your two sisters bought a gift for your brother. You paid $7.50 for your share (one-third of the gift). What was the total cost of the gift?

Answer by Theo(13342)

you paid for 1/3 of the gift.

you paid 7.50.

let x equal the price of the present.

1/3 * x = 7.50

that's your algebraic equation.

you solve the equation as follows:

1/3 * x = 7.50

multiply both sides of the equation by 3 to get:

1/3 * x * 3 = 7.50 * 3

simplify to get:

x = 22.50

that's your solution.

the price of the present is equal to x which is equal to 22.50.